# FISH-Tuning: Enhancing PEFT Methods with Fisher Information

### v1

@@ Abstract
The rapid growth in the parameter size of Large Language Models (LLMs) has led to the development of Parameter-Efficient Fine-Tuning (PEFT) methods to alleviate the computational costs of fine-tuning. Among these, Fisher Induced Sparse uncHanging (FISH) Mask is a selection-based PEFT technique that identifies a subset of pre-trained parameters for fine-tuning based on approximate Fisher information. However, the integration of FISH Mask with other PEFT methods, such as LoRA and Adapters, remains underexplored. In this paper, we propose FISH-Tuning, a novel approach that incorporates FISH Mask into addition-based and reparameterization-based PEFT methods, including LoRA, Adapters, and their variants. By leveraging Fisher information to select critical parameters within these methods, FISH-Tuning achieves superior performance without additional memory overhead or inference latency. Experimental results across various datasets and pre-trained models demonstrate that FISH-Tuning consistently outperforms the vanilla PEFT methods with the same proportion of trainable parameters.  [^1] The code for this work will be made openly accessible after the anonymous review process.
@@ 1 Introduction
FISH Mask, a selection-based PEFT method, identifies the most critical parameters for fine-tuning using Fisher information. Despite its promise, its integration with addition-based and reparameterization-based methods remains underexplored. For instance, LoRA introduces trainable low-rank weight matrices while freezing the pre-trained model weights. These matrices could also benefit from Fisher information to identify the most important parameters, potentially enhancing LoRA’s performance.
@@ 2 Related Work
2.1 Parameter Efficient Fine-tuning
With the growing size of pre-trained Large Language Models, fine-tuning all parameters is becoming increasingly expensive and may lead to overfitting Mahabadi et al. (2021). Parameter-Efficient Fine-Tuning (PEFT) has been proposed to alleviate this issue. It trains only a small proportion of the parameters while achieving similar results compared to full fine-tuning.
In Transformer architecture Vaswani et al. (2017), the Multi-Head Attention mechanism is defined as:
$\text{MultiHead}(Q,K,V)=\text{Concat}(\text{head}_{1},\dots,\text{head}_{h})W^{O}$
$\text{head}_{i}=\text{Attention}(QW_{i}^{Q},KW_{i}^{K},VW_{i}^{V})$ (2)
The feed-forward network (FFN) is given by:
$\text{FFN}(\boldsymbol{x})=\max(0,\boldsymbol{x}W_{1}+b_{1})W_{2}+b_{2}$ (3)
where $Q\in\mathbb{R}^{d},K\in\mathbb{R}^{d},V\in\mathbb{R}^{d},W_{i}^{Q}\in\mathbb{R}^{d\times k},W_{i}^{K}\in\mathbb{R}^{d\times k},W_{i}^{V}\in\mathbb{R}^{d\times k},W^{O}\in\mathbb{R}^{hk\times d},\boldsymbol{x}\in\mathbb{R}^{d},W_{1}\in\mathbb{R}^{d\times r},W_{2}\in\mathbb{R}^{r\times d},b_{1}\in\mathbb{R}^{r},b_{2}\in\mathbb{R}^{d}$. Different PEFT methods modify different weights in the Transformer. In LoRA and DoRA, $W_{Q}$, $W_{K}$, $W_{V}$, $W_{O}$, $W_{1}$, and $W_{2}$ can be used for matrix Reparameterization. The Adapter method adds an additional feed-forward layer to the Transformer. Prefix-Tuning introduces prefix matrices before $KW_{i}^{K}$ and $VW_{i}^{V}$. $(\text{IA})^{3}$ incorporates trainable vectors in $KW_{i}^{K}$, $VW_{i}^{V}$, and $\max(0,xW_{1}+b_{1})$. BitFit selects only the bias terms as trainable parameters.
@@ 3.2 FISH Mask based PEFT
$F_{\theta}=\mathbb{E}_{x\sim p(x)}[\mathbb{E}_{y\sim p_{\theta}(y|x)}\nabla_{\theta}\log p_{\theta}(y|x)\nabla_{\theta}\log p_{\theta}(y|x)^{\textrm{T}}]$
@@
$\qquad\odot\nabla_{\theta}\log p_{\theta}(y|x_{i})\Big{]}$ (5)
@@ 4 Method
We use FISH Mask into the Addition-based methods like Adapter, Prefix-Tuning, $(\text{IA})^{3}$, and the Reparameterization-based methods like LoRA, DoRA. We also use FISH Mask into the Hybrid PEFT method like UniPELT Mao et al. (2022). We believe that our method can also be used in other PEFT methods.
@@ 4.1.2 FISH Mask in DoRA
$W_{0}+\Delta W=m\frac{W_{0}+BA}{\left\|W_{0}+BA\right\|_{c}}$ (11)
@@ 4.2.1 FISH Mask in Adapter
$\text{Adapter}(\boldsymbol{x})=B\sigma(A\boldsymbol{x})+\boldsymbol{x}$ (12)
@@ 4.2.2 FISH Mask in Prefix-Tuning
$\text{head}_{i}=\text{Attention}(QW_{i}^{Q},[P_{i}^{K},KW_{i}^{K}],[P_{i}^{V},VW_{i}^{V}])$
@@ 4.2.3 FISH Mask in $(\text{IA})^{3}$
$\text{FFN}(x)=(l_{ff}\odot\sigma(Ax))B$ (15)
@@ 4.3 FISH Mask in UniPELT
$W_{0}+\Delta W=W_{0}+\mathcal{G}_{LoRA}BA$ (17)
@@ 5.1 Datasets and Baselines
Datasets. We evaluate FISH-Tuning method on the GLUE Wang et al. (2019) dataset, comparing it with the original PEFT method mentioned in Chapter 4. GLUE is a multi-task benchmark that contains 10 datasets for LLM evaluation. In our experiment, we select only the CoLA Warstadt et al. (2018), MRPC Dolan and Brockett (2005), RTE, SST-2 Socher et al. (2013), STS-B Cer et al. (2017), and WNLI Levesque et al. (2012) datasets because the remaining datasets contain too much text and require excessive training time. Then we calculate the average score of these six datasets.

### v3

@@ Abstract
The rapid growth in the parameter size of Large Language Models (LLMs) has spurred the development of Parameter-Efficient Fine-Tuning (PEFT) methods to mitigate the substantial computational costs of fine-tuning. Among these, Fisher Induced Sparse uncHanging (FISH) Mask is a selection-based PEFT technique that identifies a critical subset of pre-trained parameters using approximate Fisher information. While addition-based and reparameterization-based PEFT methods like LoRA and Adapter already fine-tune only a small number of parameters, the newly introduced parameters within these methods themselves present an opportunity for further optimization. Selectively fine-tuning only the most impactful among these new parameters could further reduce resource consumption while maintaining, or even improving, fine-tuning effectiveness. In this paper, we propose FISH-Tuning, a novel approach that incorporates FISH Mask into such PEFT methods, including LoRA, Adapter, and their variants. By leveraging Fisher information to identify and update only the most significant parameters within these added or reparameterized components, FISH-Tuning aims to achieve superior performance without increasing training time or inference latency compared to the vanilla PEFT methods. Experimental results across various datasets and pre-trained models demonstrate that FISH-Tuning consistently outperforms the vanilla PEFT methods when using the same proportion of trainable parameters. Code is available at https://anonymous.4open.science/r/FISH-Tuning-6F7C.
@@ 1 Introduction
FISH Mask is a parameter-efficient fine-tuning (PEFT) method that selects important pre-trained model parameters for fine-tuning using Fisher information. In contrast, common PEFT methods such as LoRA typically update all parameters within their added modules, like the low-rank matrices. This approach does not consider that these newly introduced parameters may also vary in importance for the downstream task. We propose that Fisher information can guide the selection of which PEFT parameters to update. This concept has received little attention so far. By applying selective fine-tuning to the internal structure of PEFT modules, it may be possible to improve both the effectiveness and efficiency of these techniques.
@@ 2 Related Work
2.1 Sparsely Updating Parameter-Efficient Fine-tuning
Parameter-Efficient Fine-Tuning (PEFT) methods aim to adapt large pre-trained language models with minimal parameter updates, and a subset of these approaches leverages sparsity to further enhance efficiency. Sparse updating in PEFT focuses on selectively modifying a small fraction of parameters or enforcing sparse structures during adaptation, reducing computational overhead and storage requirements while maintaining performance.
One prominent method in this domain is LoRI Zhang et al. (2025). Unlike conventional LoRA, which learns dense low-rank matrices $A$ and $B$, LoRI freezes $A$ as a random projection and sparsely updates only the most task-relevant entries in $B$. This design eliminates the need to store gradients and optimizer states for $A$, significantly reducing memory consumption while preserving performance through task-specific masking. Moreover, LoRI introduces an additional layer of efficiency by applying global magnitude-based sparsity to $B$, allowing up to 90% of its parameters to remain untouched during adaptation. By focusing updates on a small, critical subset of parameters, LoRI achieves strong performance while drastically lowering trainable parameter counts—up to 95% fewer than standard LoRA in some settings.
@@ 3.2 FISH Mask Based PEFT
$F_{\theta}=\mathbb{E}_{x\sim p(x)}[\mathbb{E}_{y\sim p_{\theta}(y|x)}\nabla_{\theta}\log p_{\theta}(y|x)\nabla_{\theta}\log p_{\theta}(y|x)^{\textrm{T}}]$,
@@
$\qquad\odot\nabla_{\theta}\log p_{\theta}(y|x_{i})\Big{]},$ (2)
@@ 4 Method
We use FISH Mask into the Addition-based methods like Adapter, Prefix-Tuning, $(\text{IA})^{3}$, and the Reparameterization-based methods like LoRA, DoRA. We also use FISH Mask into the Hybrid PEFT method like UniPELT Mao et al. (2022). The differences between our method and the original FISH Mask and LoRA are illustrated in Fig. 1. We believe that our method can also be used in other PEFT methods.
Figure 1: Original FISH Mask method (left) and original LoRA method (mid) and our FISH-Tuning method (right).
@@ 4.1.2 FISH Mask in DoRA
$W_{0}+\Delta W=m\frac{W_{0}+BA}{\left\|W_{0}+BA\right\|_{c}},$ (8)
@@ 4.2.1 FISH Mask in Adapter
$\text{Adapter}(\boldsymbol{x})=B\sigma(A\boldsymbol{x})+\boldsymbol{x},$ (9)
@@ 4.2.2 FISH Mask in Prefix-Tuning
$\text{head}_{i}=\text{Attention}(QW_{i}^{Q},[P_{i}^{K},KW_{i}^{K}],[P_{i}^{V},VW_{i}^{V}])$,
@@ 4.2.3 FISH Mask in $(\text{IA})^{3}$
$\text{FFN}(x)=(l_{ff}\odot\sigma(Ax))B,$ (12)
@@ 4.3 FISH Mask in UniPELT
$W_{0}+\Delta W=W_{0}+\mathcal{G}_{LoRA}BA,$ (14)
@@ 5.1 Datasets and Baselines
Datasets. We conduct experiments on two different types of tasks: classification tasks and generation tasks. For classification tasks, we select the GLUE Wang et al. (2019) dataset, and for generation tasks, we choose the GSM8K Cobbe et al. (2021) dataset. For the GLUE dataset, we use BERT Devlin et al. (2019), Modern BERT Warner et al. (2024), and LLaMA-3.2-1B Dubey et al. (2024) as the pre-trained models. For the GSM8K dataset, we use Phi-4-mini-instruct Abouelenin et al. (2025) and Qwen2.5-7B Yang et al. (2024) as the pre-trained models. The detailed introduction to these datasets and evaluation metrics will be presented in Appendix E.
We have two ways of comparison: The first method involves fixing the proportion of trainable parameters and comparing FISH-Tuning method with the original PEFT method. Since FISH-Tuning only fine-tunes a subset of parameters within specific layers, to achieve the same proportion of trainable parameters as the original PEFT method, FISH-Tuning method need to fine-tune more layers than the original PEFT method. The second method involves fixing the number of trainable layers while reducing the proportion of trainable parameters, to investigate whether there is similarity between FISH-Tuning method and the original FISH Mask method.
5.2 Implementation Details
The initialization weight method of the matrices $B$ and $A$ in LoRA and DoRA follows PiSSA Meng et al. (2024). For experiments, we did not use any hyper-parameter tuning, nor did we use MLNI trick (use the MLNI checkpoint instead of the pre-trained weights) to enhance the models’ performance. More details about the hyperparameters are available in Table 12 in Appendix D.
For the experiment fixing the proportion of trainable parameters, we select various layers of the LLMs as trainable parameters to achieve different trainable ratios. For most PEFT methods, we select the top 1, 2, 3, and 4 layers to form four groups with different trainable ratios (we consider the layers closer to the output layer as the top layers, and the layers closer to the input layer as the bottom layers). In FISH-Tuning method, we select the top 5 layers as the parameter set and use the FISH Mask to choose the top-k parameters from this set, ensuring the same trainable ratios as the original PEFT method. For $(\text{IA})^{3}$, we select the top 3, 5, 8, and 10 layers to form four groups with different trainable ratios, while in the FISH-Tuning method, we use 12 layers because $(\text{IA})^{3}$ trains fewer parameters compared to other PEFT methods.
For the experiment fixing the number of trainable layers, we select all layers of the LLMs as trainable parameters.
5.3 Experimental Results for Fixing the Proportion of Trainable Parameters
5.3.1 Baselines
Method Trainable Parameters Avg Method Trainable Parameters Avg
Original-LoRA 0.0057% 68.45 Original-PrefixTuning 0.0439% 65.40
LoRA-FISH 0.0057% 68.90 PrefixTuning-FISH 0.0439% 67.54
[…]
LoRA-FISH 0.0099% 72.25 PrefixTuning-FISH 0.0864% 67.90
[…]
LoRA-FISH 0.0142% 72.64 PrefixTuning-FISH 0.1289% 67.93
[…]
LoRA-FISH 0.0184% 72.91 PrefixTuning-FISH 0.1713% 68.49
Original-DoRA 0.0078% 68.31 Original-$(\text{IA})^{3}$ 0.0142% 65.81
DoRA-FISH 0.0078% 69.45 $(\text{IA})^{3}$-FISH 0.0142% 67.46
[…]
DoRA-FISH 0.0142% 73.09 $(\text{IA})^{3}$-FISH 0.0227% 67.99
[…]
DoRA-FISH 0.0206% 73.10 $(\text{IA})^{3}$-FISH 0.0354% 67.19
[…]
DoRA-FISH 0.0269% 73.48 $(\text{IA})^{3}$-FISH 0.0439% 68.96
Original-Adapter 0.1389% 67.85 Original-UniPELT 0.0213% 72.04
Adapter-FISH 0.1389% 73.25 UniPELT-FISH 0.0213% 69.33
[…]
Adapter-FISH 0.2760% 73.85 UniPELT-FISH 0.0411% 70.72
[…]
Adapter-FISH 0.4127% 73.82 UniPELT-FISH 0.0610% 70.89
[…]
Adapter-FISH 0.5490% 73.80 UniPELT-FISH 0.0808% 70.92
Table 1: Performance of different PEFT methods on GLUE dataset. The solid lines separate different PEFT methods, and the dashed lines separate different ratios of trainable parameters. In each dashed-line area, the first row represents the original PEFT method, and the second row represents our method. The specific details of each dataset result are represented in Appendix B Table 6.
We conduct extensive experiments on the GLUE dataset and the BERT model to verify the effectiveness of FISH-Tuning. Table 1 presents the detailed performance of FISH-Tuning and baseline methods on the GLUE benchmark. From this table, we can see that integrating the FISH Mask into various PEFT methods consistently improves performance across multiple tasks, except for UniPELT. We will discuss why FISH-Tuning performs worse in UniPELT in Limitations.
5.3.2 Results with Other Pre-trained Models
Method Trainable Parameters Avg
Original-LoRA (M) 0.0033% 63.20
LoRA-FISH 0.0033% 65.96
[…]
LoRA-FISH 0.0056% 65.75
[…]
LoRA-FISH 0.0080% 66.28
[…]
LoRA-FISH 0.0103% 66.81
Original-LoRA (L) 0.0007% 70.14
LoRA-FISH 0.0007% 71.36
[…]
LoRA-FISH 0.0010% 72.76
[…]
LoRA-FISH 0.0013% 72.96
[…]
LoRA-FISH 0.0017% 73.03
Table 2: Performance of different PEFT methods on GLUE dataset. (M) means ModernBERT-base. (L) means LLaMA-3.2-1B. The specific details of each dataset result are represented in Appendix B Table 7.
Apart from the BERT model, we also compare FISH-Tuning with the original PEFT method using other trending pre-trained models, such as ModernBERT and LLaMA-3.2-1B. The prompt template for LLaMA-3.2-1B follows Table 1 in Zhong et al. (2023). The experiment results can be seen in Table 2. These results highlight the generalizability of FISH-Tuning across different pre-trained models, confirming that FISH-Tuning is effective and robust in enhancing the performance of PEFT methods beyond the standard BERT settings.
5.3.3 Contrastive Study
Method Trainable Parameters Avg
Original-LoRA 0.0057% 68.45
LoRA-FISH 0.0057% 68.90
LoRA-FISH-rand 0.0057% 68.74
LoRA-FISH-rev 0.0057% 66.71
[…]
LoRA-FISH 0.0099% 72.25
LoRA-FISH-rand 0.0099% 69.62
LoRA-FISH-rev 0.0099% 67.93
[…]
LoRA-FISH 0.0142% 72.64
LoRA-FISH-rand 0.0142% 71.89
LoRA-FISH-rev 0.0142% 71.86
[…]
LoRA-FISH 0.0184% 72.91
LoRA-FISH-rand 0.0184% 72.92
LoRA-FISH-rev 0.0184% 72.65
Table 3: Performance of different PEFT methods on GLUE dataset. In each dashed-line area, the first row represents the original method, the second row represents our method, the third row represents the method where we randomly select the important parameters without using the FISH Mask, and the fourth row represents the method where we select the important parameters in reverse order compared to the second method. The specific details of each dataset result are represented in Appendix B Table 8.
We further investigate the impact of the FISH Mask by conducting a contrastive study on the LoRA framework, as detailed in Table 3. In this study, we compare the standard LoRA method with three variants: our proposed LoRA-FISH (the standard FISH-Tuning method), a variant where important parameters are selected at random (LoRA-FISH-rand), and another variant where the importance ordering is reversed (LoRA-FISH-rev). The parameter selection method of LoRA-FISH-rev is described as follows, replacing Equation 4:
The results from 3 clearly indicate that the deliberate selection of important parameters using the FISH Mask is key to achieving the performance gains.
5.4 Experimental Results for Fixing the Number of Trainable Layers
5.4.1 Result on Classification Task
Figure 4: Original FISH Mask method (left) and our FISH-Tuning method (right). The specific details of each dataset result of the right picture are represented in Appendix B Table 9.
Fig. 4 presents the detailed performance of FISH-Tuning and baseline methods on the GLUE benchmark. In Fig. 4, the experimental results on the left side of the image are from Sung et al. (2021), and the results on the right side are from our experiments. The "proportion of trainable parameters" on the left side represents the ratio of trainable parameters to the total model parameters, whereas the "proportion of trainable parameters" on the right side refers to the ratio of trainable parameters to the total parameters within the LoRA component. Although FISH Tuning selects the most important subset of parameters from randomly initialized parameters, whereas the original FISH Mask selects the most important subset from the pre-trained model’s parameters, both approaches yield similar patterns: (1) The parameter selection method of FISH Mask outperforms both Random Mask and Reverse Mask. (2) When the proportion value is relatively large, the performance differences among Random Mask, Reverse Mask, and FISH Mask become smaller.
5.4.2 Result on Generation Task
Figure 5: Performance of FISH-Tuning method in different proportion of Trainable Parameters on the GSM8K dataset using Qwen2.5-7B model after 4 training epochs.
We also conduct experiments on the generation task using the Qwen2.5-7B model. Fig. 5 displays the performance results of FISH-Tuning on the GSM8K dataset after completing 4 training epochs. As shown in Fig. 5, when the proportion value is small, the result of FISH Mask is significantly higher than that of Random Mask, and the result of Random Mask is clearly higher than that of Reverse Mask. Unlike Fig. 4, as the proportion value gradually decreases, the performance of Random Mask and Reverse Mask drops sharply, while the result of FISH Mask hardly declines at all—sometimes even surpassing Dense fine-tuning. This phenomenon indicates that, when training generative tasks with a very small number of trainable parameters, selecting parameters for fine-tuning based on Fisher Information can effectively prevent deterioration in model training outcomes. Pictures for other training epochs are included in the Appendix F, Fig. 8.
Figure 6: LoRA-FISH converges faster compared to LoRA-FISH-random and LoRA-FISH-reverse.
Fig. 6 shows the performance of FISH-Tuning across different training epochs when the proportion value is set to 1%. From Fig. 6, we observe that FISH Mask nearly converges after just 1 training epoch, while Random Mask approaches convergence only after 7 epochs. In contrast, Reverse Mask maintains an accuracy rate of zero throughout. By the 10-th training epoch, the result of FISH Mask is significantly better than that of Random Mask, and Random Mask is significantly better than Reverse Mask. The results presented in Fig. 6 further confirm the importance of using Fisher Information to select parameters for fine-tuning. Pictures for other proportion values are provided in the Appendix F, Fig. 9.
In addition to Qwen2.5-7B, similar experiments were conducted on Phi-4-mini-instruct, where comparable results are observed. The experimental results for Phi-4-mini-instruct are detailed in the Appendix G, Fig. 10 and Fig. 11.
5.4.3 Comparison with LoRI
Method Trainable parameters in LoRA (‰)
0.018 0.036 0.18 0.36 0.9
LoRI 78.20 79.14 79.53 78.13 77.50
FISHTuning 78.24 76.88 78.01 78.77 78.47
Table 4: The performance of LoRI and our FISH-Tuning method in the same proportion of trainable parameters on Qwen2.5-7B.
We compared our method with the LoRI Zhang et al. (2025) approach, and the comparison results are shown in Table 4. We set the same hyperparameters to ensure fairness. As illustrated in Table 4, our FISH-Tuning method outperforms LoRI under three proportions of trainable parameter ratios, while it performs worse than LoRI under two proportions. Overall, these experimental results validate the effectiveness of FISH-Tuning in balancing model performance and parameter selection, demonstrating its potential as a novel fine-tuning approach.
5.4.4 Resource Consumption Analysis
We compare the changes in resource consumption caused by adding FISH-Tuning to the original PEFT methods. Table 5 presents the resource consumption when training a BERT model on the RTE task for 1 epoch, under the condition of fixing the number of trainable layers. The first row in Table 5 shows the runtime and GPU memory usage when using the original LoRA method for training, and the other rows show the runtime and GPU memory usage when fine-tuning part of the parameters in LoRA component after using FISH-Tuning. As the proportion of trainable parameters decreases, the required runtime decreases while GPU memory slightly increases. The runtime reduction results from fewer parameters requiring fine-tuning within the model, while the increase in GPU memory is due to additional matrices introduced by the FISH Mask consuming memory. In theory, if GPUs were to provide hardware-level support for sparse matrix parameter updates, the GPU memory consumption would also decrease. Incorporating FISH-Tuning does not affect the time or GPU memory consumed during inference since the FISH Mask is only used during training.
Method Ratio1 Ratio2 Time GPU
Origin-LoRA 0.8116% 100% 38.0208 6180
LoRA-FISH 0.4058% 50% 37.6212 6650
LoRA-FISH 0.2029% 25% 37.4502 6650
LoRA-FISH 0.0812% 10% 37.7291 6650
LoRA-FISH 0.0406% 5% 37.0493 6650
LoRA-FISH 0.0081% 1% 37.1929 6650
LoRA-FISH 0.0041% 0.5% 36.6827 6650
Table 5: Resource consumption of FISH-Tuning method in different proportion of Trainable Parameters. The "Ratio1" column represents the ratio of trainable parameters to the total number of parameters, the "Ratio2" column represents the ratio of trainable parameters to the total number of parameters in LoRA component, the "Time" column represents the training time in seconds, and the "GPU" column represents the GPU memory consumption in MB.
6 Conclusion
In this paper, we propose FISH-Tuning, a new method that uses the selective PEFT method in the Addition-based and Reparameterization-based PEFT methods. More concretely, we integrate the FISH Mask into LoRA, DoRA, Adapter, Prefix-Tuning, and $(\text{IA})^{3}$. With the same ratio of trainable parameters, our method outperforms the original PEFT method most of the time. With the same number of trainable layers and ranks, our method achieves experimental results similar to those of the original FISH Mask method. In future work, we will explore our approach in Computer Vision, Multi-modality, and Quantization.
Limitations
2) We use the LoRA method as an example. To achieve the same ratio of trainable parameters in FISH-Tuning, we can select different trainable layers and LoRA ranks. Tables 10 and 11 in Appendix C demonstrate our experiments on this. According to the results, we have not found any obvious pattern for determining how to select the suitable ranks and layers to achieve the best result.
Nevertheless, we are excited to see the huge potential of FISH-Tuning already demonstrated in existing experiments and look forward to more tests and suggestions from the community. In future work, we will try to incorporate other Selection-based PEFT methods into Addition-based PEFT methods and Reparameterization-based PEFT methods and compare them with our FISH-Tuning method.
Ethics Statement
Acknowledgments
Appendix A Loss on Evaluation Datasets
Figure 7: Original LoRA method loss (Blue line) and the LoRA-FISH method loss (Orange line). The x-axis represents the training epochs, and the y-axis denotes the average evaluation loss scores across six datasets.
We conduct an experiment on the loss scores using the evaluation dataset. The loss values are taken from the first solid-line area in Table 1. We visualize the loss value curves for the original PEFT method and FISH-Tuning method in Fig. 7. From this figure, we observe that all points on the FISH-Tuning curve are lower than those on the original PEFT method curve. This result demonstrates that the FISH-Tuning method converges faster than the original PEFT method. It also explains why FISH-Tuning method achieves better results in Table 1.
Appendix B Tables of Precision Results
This section will present the model performance results on the GLUE dataset mentioned in the main text. Due to space limitations, the tables in the main text only list the average scores of each model on the GLUE dataset, whereas this section will further provide the detailed performance metrics of each model on individual subtasks of GLUE.
Table 6 and Table 7 present the detailed performance of FISH-Tuning and base-line methods on the GLUE benchmark. The pre-trained model of Table 6 is bert-base-cased. The pre-trained models of Table 7 are ModernBERT-base and LLaMA-3.2-1B.
Table 8 compare the standard LoRA method with three variants: our proposed LoRA-FISH (the standard FISH-Tuning method), a variant where important parameters are selected at random(LoRA-FISH-rand), and another variant where the importance ordering is reversed (LoRA-FISH-rev). The pre-trained model is bert-base-cased.